Required information

[The following information applies to the questions displayed below.]

Westerville Company reported the following results from last year’s operations:

Sales $ 1,000,000
Variable expenses 300,000
Contribution margin 700,000
Fixed expenses 500,000
Net operating income $ 200,000
Average operating assets $ 625,000


At the beginning of this year, the company has a $120,000 investment opportunity with the following cost and revenue characteristics:

Sales $ 200,000
Contribution margin ratio 60 % of sales
Fixed expenses $ 90,000

The company’s minimum required rate of return is 15%.

2. What is last year’s turnover? (Round your answer to 1 decimal place.)

3. What is last year’s return on investment (ROI)?

11. What is last year’s residual income?
Answer #1

2) Turnover = Sales/Average operating assets = 1000000/625000 = 1.6

3) Return on investment = Net operating income/Average operating assets = 200000/625000 = 32%

11) Residual income = 200000-(625000*15%) = 106250